# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 6.3 Projected Primal Polytope Approximates Two-Dimensional Unit Disk
Lemma 51 (Projected primal polytope approximates the two-dimensional disk).
@@
Lemma 51 directly follows from the next two lemmas. First, we show that the two-dimensional unit disk is contained in $\pi_{W}(P)$:
Lemma 52 (Inner radius of the projected primal polytope).
@@
Proof.
Given such $x,y$, set $s=0$ and set $p_{0},p_{1},\cdots,p_{k},t$ such that (22) and (24) are satisfied with equality. This will result in $\sqrt{x^{2}+y^{2}}=\|p_{0}\|=\|p_{i}\|$ for every $i\in[k]$. Since $t=|v_{i}^{\top}p_{i-1}|\leq\|v_{i}\|\cdot\|p_{i-1}\|\leq 1$, we know that (26) is satisfied. Furthermore, we have $e_{1}^{\top}p_{k}\leq\|e_{1}\|\cdot\|p_{k}\|=\|e_{1}\|\cdot\sqrt{x^{2}+y^{2}}\leq 1$ which ensures that (25) is satisfied. ∎
In the next lemma, we will show that $\pi_{W}(P)$ is contained in the two-dimensional disk $\cos(\pi/2^{k+2})^{-1}\mathbb{B}_{2}^{2}$:
Lemma 53 (Outer radius of the projected primal polytope).
@@
Proof.
Fix any $(x,y)\in\mathbb{R}^{2}$ and $p_{0}\in\mathbb{R}^{2}$, such that $\sqrt{x^{2}+y^{2}}>\cos(\pi/2^{k+2})^{-1}$ and $p_{0}\geq\begin{bmatrix}|x|,|y|\end{bmatrix}^{\top}$. Also fix any $p_{1},\ldots,p_{k}\in\mathbb{R}^{2}$ satisfying (23) and (24). We will show that such $p_{1},\ldots,p_{k}$ would violate (25), i.e. $e_{1}^{\top}p_{k}>1$. To simplify our notation, for all $i\in\{0,1,\ldots,k\}$, let $(p_{i})_{v}=v_{i}^{\top}p_{i}\in\mathbb{R}$ and $(p_{i})_{w}=w_{i}^{\top}p_{i}\in\mathbb{R}$. Then $p_{i}=(p_{i})_{v}v_{i}+(p_{i})_{w}w_{i}$.
@@
Thus we have shown $x^{2}+y^{2}>\cos(\pi/2^{k+2})^{-1}$ implies that $e_{1}^{\top}p_{k}\geq e_{1}^{\top}p_{k}^{*}>1$ as desired. ∎
6.4 Inner and Outer Radius of the Primal Polytope
In this subsection, we will show that the primal polytope $P$ has large inner radius and small outer radius.
Lemma 54 (Inner radius of the primal polytope).
@@ 6.4 Inner and Outer Radius of the Primal Polytope
Proof.
In Lemma 55 we construct a point $(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$ such that $\frac{1}{30}\mathbb{B}_{\infty}^{k+5}\subseteq P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$.
For the second inclusion, we show that $P\subseteq(\cos(\pi/2^{k+2})^{-1}\cdot\mathbb{B}_{\infty}^{k+5}$. Suppose $(x,y,p_{0},t,s)\in P$ is arbitrary. From Lemma 53 we know that $\|(x,y)\|_{\infty}\leq\sqrt{x^{2}+y^{2}}\leq\|p_{0}\|_{2}\leq\cos(\pi/2^{k+2})^{-1}$. Since $\mathbf{0}_{k}\leq t\leq\mathbf{1}_{k}$ we get $\|t\|_{\infty}\leq 1$, and lastly we have $0\leq s\leq 1$. Put together, we find that $\|(x,y,p_{0},t,s)\|_{\infty}\leq\cos(\pi/2^{k+2})^{-1}$. Since $(x,y,p_{0},t,s)\in P$ was arbitrary, we find that $P\subseteq\cos(\pi/2^{k+2})^{-1}\cdot\mathbb{B}_{\infty}^{k+5}$. By the triangle inequality we find that $P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\subseteq(\cos(\pi/2^{k+2})^{-1}+\|(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\|_{\infty})\cdot\mathbb{B}_{\infty}^{k+5}$ and we will see in Lemma 55 that $\|(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\|_{\infty}=1/3$. ∎
Lemma 55 (Inner Radius of the Polytope).
@@
Proof.
Fix any $(x,y,p_{0},t,s)\in\mathbb{R}^{k+5}$ such that $\|(x-\bar{x},y-\bar{y},p_{0}-\bar{p}_{0},t-\bar{t},s-\bar{s})\|_{\infty}\leq r$. Let $p_{1},\ldots,p_{k}\in\mathbb{R}^{2}$ be uniquely defined by (23) and the equality in (24), i.e., $w_{i}^{\top}p_{i}=w_{i}^{\top}p_{i-1}$ and $t_{i}+is=v_{i}^{\top}p_{i}$ with our fixed $p_{0},s$ and $t$. We will show that $(x,y,p_{0},t,s)\in P$ by verifying (22 - 27). To simplify our notation, we let $(p_{i})_{v}=v_{i}^{\top}p_{i}\in\mathbb{R}$ and $(p_{i})_{w}=w_{i}^{\top}p_{i}\in\mathbb{R}$ for all $i\in\{0,1,\ldots,k\}$. Then $p_{i}=(p_{i})_{v}v_{i}+(p_{i})_{w}w_{i}$.
First, observe that $e_{1}^{\top}p_{0}\geq\frac{1}{6}-r\geq r\geq|x|$ and $e_{2}^{\top}p_{0}\geq\frac{1}{6}-r\geq r\geq|y|$, confirming that (22) holds. Also, notice that $t_{i}\in[\bar{t}_{i}-r,\bar{t_{i}}+r]\subseteq[0,1]$ and $s\in[\bar{s}-r,\bar{s}+r]\subseteq[0,1]$. Thus (26) and (27) hold. The equality constraint (23) holds by definition of $p_{1},\ldots,p_{k}$.
It remains to show (24) and (25), i.e. $(p_{i})_{v}\geq|v_{i}^{\top}p_{i-1}|$ for all $i\in[k]$ and $e_{1}^{\top}p_{k}\leq 1$. To aid in the remaining steps of the proof, we show $(p_{i})_{w}\geq 0$ for all $i\in\{0,1,\ldots,k\}$. Observe that $w_{0}^{\top}p_{0}\geq w_{0}^{\top}\bar{p}_{0}-\|w_{0}\|\cdot\|p_{0}-\bar{p}_{0}\|\geq\frac{\sqrt{2}}{6}-\sqrt{2}r\geq 0$. Also, for all $i\in[k]$,
@@
$\geq(p_{i-1})_{w}\cdot\cos(\frac{\pi}{2^{i+2}}).$
It then follows by induction that $(p_{i})_{w}\geq 0$ for all $i\in\{0,1,\ldots,k\}$.
Next, we verify the inequality listed in (24), i.e., $(p_{i})_{v}\geq|v_{i}^{\top}p_{i-1}|$ for all $i\in[k]$. Notice that for all $i\in[k]$,
@@
Let $t_{0}=v_{0}^{\top}p_{0}$ and $\bar{t}_{0}=v_{0}^{\top}\bar{p}_{0}=0$. By applying the above inequality to $11,2,\cdots,i-$ 1, we have
@@
Finally, we verify (25). The increment of the first coordinate from $p_{i-1}$ to $p_{i}$ is
@@
where the last step comes from $v_{i}^{\top}v_{i-1}=\cos(\frac{\pi}{2^{i+2}})$ and $v_{i}^{\top}w_{i-1}=-\sin(\frac{\pi}{2^{i+2}})$. For all $i\geq 2$, we can show that in (38), the third term in the brackets is at most the fourth term:
@@
$\leq(0.338+0.898\cdot\frac{11}{30})\cdot\tan(\frac{\pi}{8})\cdot\cos(\frac{\pi}{2^{i+2}})$ (By $r\leq\frac{1}{30}$ and $s\leq\bar{s}+r=\frac{11}{30}$)
$\leq 0.277\cdot\cos(\frac{\pi}{2^{i+2}})$
$\leq(\bar{t}_{i-1}-r+(i-1)s)\cdot\cos(\frac{\pi}{2^{i+2}})$ (By $\bar{t}_{i-1}-r=0$ and $s\geq\bar{s}-r=0.3$)
@@
Plugging back into (38), we have
@@
Therefore,
@@
$\leq(\frac{1}{6}+r)+\sin(\frac{\pi}{8})\cdot\sum_{i=1}^{k}1.9^{-(i-1)}\cdot(\frac{1}{30}+\frac{i}{3}+(i+1)r)$
$\leq 0.763+3.514r\leq 1.$
where the last inequality holds for any $r\leq\frac{1}{30}$. Therefore (25) holds and $(x,y,p_{0},t,s)\in P$. ∎
6.5 Properties of the Dual Polytope
In this section, we will analyze the scaled polar dual polytope $Q=\frac{1}{30}(P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s}))^{\circ}$. From well-known duality properties, we will find that $Q$ satisfies the following desirable properties:

### v3

@@ 6.2 Projected Auxiliary Polytope Approximates Two-Dimensional Unit Disk
Lemma 6.3 (Projected auxiliary polytope approximates the two-dimensional disk).
@@
Lemma 6.3 directly follows from the next two lemmas. First, we show that the two-dimensional unit disk is contained in $\pi_{W}(P)$:
Lemma 6.4 (Inner $\ell_{2}$-radius of the projected auxiliary polytope).
@@
Proof 6.5.
Suppose $x,y\in\mathbb{R}$ with $x^{2}+y^{2}\leq 1$. We want to exhibit $p_{0}\in\mathbb{R}^{2}$, $t\in\mathbb{R}^{k}$, and $s\in\mathbb{R}$ such that $(x,y,p_{0},t,s)\in P$. We proceed as follows:
First, set $s=0$ and define $p_{0}=(x,y)$. Then define $p_{i}$ inductively by the recurrence
$p_{i}=\begin{bmatrix}v_{i}^{\top}\\ w_{i}^{\top}\end{bmatrix}^{-1}\begin{bmatrix}\lvert v_{i-1}^{\top}p_{i-1}\rvert\\ w_{i-1}^{\top}p_{i-1}\end{bmatrix}$,
with the base case being $p_{0}=(|x|,|y|)$. Finally, for each $i\in[k]$ define $t_{i}=v_{i}^{\top}p_{i}=|v_{i}^{\top}p_{i-1}|$.
Then (24), (25), (26) and (29) follow by definition. It remains to argue that (27) and (28) are satisfied. Since $v_{i}$ and $w_{i}$ are orthogonal and defined to have norm $1$, the matrix $\begin{bmatrix}v_{i}^{\top}\\ w_{i}^{\top}\end{bmatrix}$ is an isometry. Thus, it preserves norms, meaning that each of the $p_{i}$ have $\ell_{2}$-norm $\sqrt{x^{2}+y^{2}}$. Since $t_{i}=|v_{i}^{\top}p_{i-1}|\leq\|v_{i}\|\cdot\|p_{i-1}\|\leq 1$, we know that (28) is satisfied. Furthermore, we have $e_{1}^{\top}p_{k}\leq\|e_{1}\|\cdot\|p_{k}\|=\|e_{1}\|\cdot\sqrt{x^{2}+y^{2}}\leq 1$, which ensures that (27) is satisfied.
In the next lemma, we show that the projection $\pi_{W}(P)$ is contained in the two-dimensional disk $\cos(\pi/2^{k+2})^{-1}\mathbb{B}_{2}^{2}$:
Lemma 6.6 (Outer $\ell_{2}$-radius of the projected auxiliary polytope).
@@
Proof 6.7.
Fix any $(x,y)\in\mathbb{R}^{2}$ and $p_{0}\in\mathbb{R}^{2}$, such that $\sqrt{x^{2}+y^{2}}>\cos(\pi/2^{k+2})^{-1}$ and $p_{0}\geq\begin{bmatrix}|x|,|y|\end{bmatrix}^{\top}$. Also fix any $p_{1},\ldots,p_{k}\in\mathbb{R}^{2}$ satisfying (25) and (26). We will show that such $p_{1},\ldots,p_{k}$ would violate (27), i.e., $e_{1}^{\top}p_{k}>1$. To simplify our notation, for all $i\in\{0,1,\ldots,k\}$, let $(p_{i})_{v}=v_{i}^{\top}p_{i}\in\mathbb{R}$ and $(p_{i})_{w}=w_{i}^{\top}p_{i}\in\mathbb{R}$. Then $p_{i}=(p_{i})_{v}v_{i}+(p_{i})_{w}w_{i}$.
@@
Thus we have shown $x^{2}+y^{2}>\cos(\pi/2^{k+2})^{-1}$ implies that $e_{1}^{\top}p_{k}\geq e_{1}^{\top}p_{k}^{*}>1$ as desired.
6.3 Inner and Outer $\ell_{\infty}$-Radius of the Auxiliary Polytope
In this subsection, we will show that the auxiliary polytope $P$ has large inner $\ell_{\infty}$-radius and small outer $\ell_{\infty}$-radius.
Lemma 6.8 (Inner and Outer $\ell_{\infty}$-Radius of the Auxiliary Polytope).
@@ 6.3 Inner and Outer $\ell_{\infty}$-Radius of the Auxiliary Polytope
The following lemma is the key to show Lemma 6.8, where we construct a point in $B$ such that the $\ell_{\infty}$ ball with radius $\frac{1}{30}$ centered at that point is contained in $P$. Obtaining a large inner ball is relatively easy using linear programming. In this exposition we demonstrate the large inner ball by hand, which will require a somewhat tedious calculation to 3 significant digits.
Lemma 6.9 (Inner $\ell_{\infty}$-Radius of the Polytope).
@@
Proof 6.10.
Fix any $(x,y,p_{0},t,s)\in\mathbb{R}^{k+5}$ such that $\|(x-\bar{x},y-\bar{y},p_{0}-\bar{p}_{0},t-\bar{t},s-\bar{s})\|_{\infty}\leq r$. We set $p_{1},\ldots,p_{k}\in\mathbb{R}^{2}$ inductively by (25) and the equations in (26), i.e., $w_{i}^{\top}p_{i}=w_{i}^{\top}p_{i-1}$ and $t_{i}+is=v_{i}^{\top}p_{i}$ start from the base case of $p_{0},s$ and $t$.
We will show that $(x,y,p_{0},t,s)\in P$ by verifying the conditions in (24 - 29). To simplify our notation, we define $(p_{i})_{v}=v_{i}^{\top}p_{i}\in\mathbb{R}$ and $(p_{i})_{w}=w_{i}^{\top}p_{i}\in\mathbb{R}$ for all $i\in\{0,1,\ldots,k\}$. Note that $v_{i}$ and $w_{i}$ are orthogonal, $p_{i}=(p_{i})_{v}v_{i}+(p_{i})_{w}w_{i}$.
First, observe that $e_{1}^{\top}p_{0}\geq\frac{1}{6}-r\geq r\geq|x|$ and $e_{2}^{\top}p_{0}\geq\frac{1}{6}-r\geq r\geq|y|$, confirming that (24) holds. Also, notice that $t_{i}\in[\bar{t}_{i}-r,\bar{t_{i}}+r]\subseteq[0,1]$ and $s\in[\bar{s}-r,\bar{s}+r]\subseteq[0,1]$. Thus (28) and (29) hold. The equality constraint (25) holds directly by definition of $p_{1},\ldots,p_{k}$. It remains to show (26) and (27), i.e., $(p_{i})_{v}\geq|v_{i}^{\top}p_{i-1}|$ for all $i\in[k]$ and $e_{1}^{\top}p_{k}\leq 1$.
To aid in the remaining steps of the proof, we show the claim that $(p_{i})_{w}\geq 0$ for all $i\in\{0,1,\ldots,k\}$ by induction. Observe that $w_{0}^{\top}p_{0}\geq w_{0}^{\top}\bar{p}_{0}-\|w_{0}\|\cdot\|p_{0}-\bar{p}_{0}\|\geq\frac{\sqrt{2}}{6}-\sqrt{2}r\geq 0$. Also, for all $i\in[k]$,
@@
$\geq(p_{i-1})_{w}\cdot\cos(\frac{\pi}{2^{i+2}}),$
where (32) follows from $w_{i}^{\top}w_{i-1}=\|w_{i}\|\|w_{i-1}\|\cos(\theta)=\cos(\theta)$ where $\theta=\pi/2^{i+2}$ is the angle between $w_{i}$ and $w_{i-1}$; similarly, we have $w_{i}^{\top}v_{i-1}=\|w_{i}\|\|v_{i-1}\|\cos(\pi/2-\theta)=\sin(\theta)=\sin(\pi/2^{i+2})$. It then follows by induction that $(p_{i})_{w}\geq 0$ for all $i\in\{0,1,\ldots,k\}$.
Verify (26):
To verify the inequality listed in (26), i.e., $(p_{i})_{v}\geq|v_{i}^{\top}p_{i-1}|$ for all $i\in[k]$, we upper-bound the right-hand-side by expanding it into the inner product with $v_{i-1}$ and the inner product with $w_{i-1}$. Notice that for all $i\in[k]$,
@@
Let $t_{0}=v_{0}^{\top}p_{0}$ and $\bar{t}_{0}=v_{0}^{\top}\bar{p}_{0}=0$. By applying the above inequality to $1,2,\cdots,i-1$, we have
@@
$\leq\frac{\sqrt{2}}{6}+\sin(\frac{\pi}{8})\cdot\frac{1}{30}\cdot\sum_{j=0}^{\infty}1.9^{-j}+r\left(\sqrt{2}+\sin(\frac{\pi}{8})\sum_{j=0}^{\infty}1.9^{-j}\right)+s\left(\sin(\frac{\pi}{8})\sum_{j=0}^{\infty}j(1.9)^{-j}\right)$
@@
Here the second inequality uses the fact that for every $x\in[0,\frac{\pi}{8}]$ one has $\sin(x)/1.9\geq\sin(x/2)$. This is because by the half angle formula, $\sin(x/2)=\pm\sqrt{\frac{1-\cos(x)}{2}}$, thus
$\frac{\sin(x)^{2}}{\sin(x/2)^{2}}=\frac{1-\cos(x)^{2}}{(1-\cos(x))/2}=2+2\cos(x)\geq 2+2\cos(\pi/8)\geq 1.9^{2}.$
@@
Verify (27):
To verify (27), notice that increment of the first coordinate from $p_{i-1}$ to $p_{i}$ is
@@
where the last step comes from $v_{i}^{\top}v_{i-1}=\cos(\frac{\pi}{2^{i+2}})$ and $v_{i}^{\top}w_{i-1}=-\sin(\frac{\pi}{2^{i+2}})$. For all $i\geq 2$, we can show that in (35), the third term in the brackets is at most the fourth term, thus the right-hand-side of (35) is at most $\sin(\pi/2^{i+2})\cdot(t_{i}+is)$:
@@
$\leq 0.277\cdot\cos(\frac{\pi}{2^{i+2}})$ (By $r\leq\frac{1}{30}$ and $s\leq\bar{s}+r=\frac{11}{30}$)
$\leq(\bar{t}_{i-1}-r+(i-1)s)\cdot\cos(\frac{\pi}{2^{i+2}})$ (By $\bar{t}_{i-1}-r=0$ and $(i-1)s\geq s\geq\bar{s}-r=0.3$)
@@
Plugging back into (35), we have for all $i\geq 2$ that
@@
For $i=1$ we recall from above that $(p_{i-1})_{w}\cdot\sin(\frac{\pi}{2^{i+2}})\leq 0.277\cdot\cos(\frac{\pi}{2^{i+2}})$. Now take (35) and observe $e_{1}^{\top}p_{1}-e_{1}^{\top}p_{0}=\sin(\pi/8)(t_{1}+s+(p_{0})_{w}\cdot\sin(\pi/8)-t_{0}\cdot\cos(\pi/8))\leq\sin(\pi/8)(11/30+2r+0.277\cdot\cos(\pi/8))$. Hence we find that $e_{1}^{\top}p_{1}-e_{1}^{\top}p_{0}\leq 0.239+0.766r$ and therefore,
@@
$\leq(\frac{1}{6}+r)+0.239+0.766r+\sin(\frac{\pi}{8})\cdot\sum_{i=2}^{k}1.9^{-(i-1)}\cdot(\frac{1}{30}+\frac{i}{3}+(i+1)r)$
$\leq(\frac{1}{6}+r)+0.239+0.766r+0.456+1.749r$
$\leq 0.862+3.515r\leq 1.$
where the last inequality holds for any $r\leq\frac{1}{30}$. Therefore (27) holds and $(x,y,p_{0},t,s)\in P$.
Proof 6.11 (Proof of Lemma 6.8).
In Lemma 6.9 we construct a point $(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$ such that $\frac{1}{30}\mathbb{B}_{\infty}^{k+5}\subseteq P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})$.
For the second inclusion, we show that $P\subseteq(\cos(\pi/2^{k+2})^{-1}\cdot\mathbb{B}_{\infty}^{k+5}$. Suppose $(x,y,p_{0},t,s)\in P$ is arbitrary. From Lemma 6.6 we know that $\|(x,y)\|_{\infty}\leq\sqrt{x^{2}+y^{2}}\leq\|p_{0}\|_{2}\leq\cos(\pi/2^{k+2})^{-1}$. Since $\mathbf{0}_{k}\leq t\leq\mathbf{1}_{k}$ we get $\|t\|_{\infty}\leq 1$, and lastly we have $0\leq s\leq 1$. Put together, we find that $\|(x,y,p_{0},t,s)\|_{\infty}\leq\cos(\pi/2^{k+2})^{-1}$. Since $(x,y,p_{0},t,s)\in P$ was arbitrary, we find that $P\subseteq\cos(\pi/2^{k+2})^{-1}\cdot\mathbb{B}_{\infty}^{k+5}$. By the triangle inequality we find that $P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\subseteq(\cos(\pi/2^{k+2})^{-1}+\|(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\|_{\infty})\cdot\mathbb{B}_{\infty}^{k+5}$ and we see in Lemma 6.9 that $\|(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s})\|_{\infty}=1/3$. Finally, note that $\cos(\pi/2^{k+2})^{-1}\leq\cos(\pi/8)^{-1}\leq 1.1$, we have
$\cos(\frac{\pi}{2^{k+2}})^{-1}\leq\cos(\frac{\pi}{8})^{-1}+\frac{1}{3}\leq\frac{3}{2}.$
6.4 Properties of the Polar Polytope
In this section, we will analyse the scaled polar polytope $Q=\frac{1}{30}(P-(\bar{x},\bar{y},\bar{p}_{0},\bar{t},\bar{s}))^{\circ}$. From well-known duality properties, we will find that $Q$ satisfies the following desirable properties: